OpenStudy (anonymous):

Night engineers are going on an energy assessment in 3 cars that hold 2, 3, and 4 passengers, respectively. In how many ways is it possible to transport the 9 engineers to the manufacturing facility, using all cars?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (kropot72):

There are 9!/2!7! combinations of 2 engineers for the car that holds 2 passengers. Having selected a combination of engineers for the car that holds 2 passengers, there are 7!\3!4! combinations of 3 engineers for the car that holds 3 passengers. Having selected combinations of engineers for the cars that hold 2 and 3 passengers there are only 4 engineers for the car that holds 4 passengers, giving only one combination. Each combination of 2 engineers can be taken with each of the combinations of 3 engineers. So the total number of ways it is possible to transport the engineers is \[\frac{9!}{2!7!}\times\frac{7!}{3!4!}\times1=\frac{9\times8\times7\times6\times5}{2\times3\times2}=you\ can\ calculate\]